OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2

OpenStudy (anonymous):

point slope form

OpenStudy (anonymous):

y+2=2(x-3)

OpenStudy (anonymous):

y-(-2)=2(x-3) y=2x-8

OpenStudy (anonymous):

\[y-y0=m(x-x0)\]

OpenStudy (anonymous):

Plug in x0, y0 and m values and get the reqd eqn.